OpenStudy (anonymous):

What is the equation of the line that passes through the point (1, 6) and has a slope of 3?

OpenStudy (anonymous):

y = 3x + 5 y = 3x + 3 y = 3x - 17 y = 3x - 5

OpenStudy (amistre64):

y = mx-mPx+Py

OpenStudy (amistre64):

y = 3x -3(1)+6

OpenStudy (nenadmatematika):

y=3x+3

OpenStudy (anonymous):

y = mx + b 6 = 3*1 + b => b = 6 - 3 = 3 Thus y = 3x + 3